Valerie has a bag of marbles that contains 8 red,4 green, 3 orange,5 blue, and 4 purple marbles. If she pulls out 2 marbles at random, one at a time without replacing the chosen marbles,then what is the probability that both are purple?
Answer by Nate(3500)

8-red
4-green
3-orange
5-blue
4-purple
Sum of Balls = 24
Chance of Purple * Second Chance of Purple
(4/24)(3/23) = (1/6)(3/23) = 3/138 = 0.02173913 or about 2.2%